Two cars leave from Port of Spain and travel s west. One travels 1 1/2 times as fast as the other. After 3 hours they are 45 miles apart. What are the speeds of the cars?
Answer by checkley77(12844)

D=RT
D=(1.5X-X)3
45=.5X*3
45=1.5X
1.5X=45
X=45/1.5
X=30 ANS.
PROOF:
45=(1.5*30-30)3
45=(45-30)3
45=15*3
45=45